OpenStudy (anonymous):

Find the derivative of f(x) = 6/x at x = -2

OpenStudy (anonymous):

x=-3

OpenStudy (anonymous):

easy

OpenStudy (anonymous):

take the derivative first algebra

OpenStudy (anonymous):

-6/x^2 set x=2

OpenStudy (anonymous):

thats not it

OpenStudy (ddcamp):

f(x) = 6(x^-1) f'(x) = -6(x^-2) -2^-2 = 1/4 → f'(-2) = -1.5

OpenStudy (anonymous):

tholn did u try my method?

OpenStudy (anonymous):

these are the choices they give us 6 3/2 -3/2 -6

OpenStudy (ddcamp):

-1.5 = -3/2

OpenStudy (anonymous):

oh ok i like your method

OpenStudy (anonymous):

i agree with champ

OpenStudy (anonymous):

i have another one i need help with